OpenStudy (anonymous):

Solve. (x – 5)2 = 4. x = 3 and x = 11 x = 9 and x = 11 x = 9 and x = 1 x = 7 and x = 3

OpenStudy (anonymous):

Don't give answer directly @imnotamoose

OpenStudy (anonymous):

Delete that one..

OpenStudy (karatechopper):

Ok well lets just starrt solving for X. Our equation is: (x – 5)2 = 4 First we will distribute the 2 in this. So... (x – 5)2 = 4 will become.. 2x-10=4 Then we will add the 10 to both sides. 2x-10=4 +10+10 ---------- 2x=14 Next, we have to divide the 2 on both sides.. 2x=14 2x/2=14/2 x=7 Now look at the choices given... x = 3 and x = 11 x = 9 and x = 11 x = 9 and x = 1 x = 7 and x = 3 There is only one answer choice with x=7. Therefore the last answer choice is your answer.

OpenStudy (karatechopper):

But its against the CoC to give the answer:/. This is a helping site!! not answer directly!

OpenStudy (anonymous):

Please read: http://openstudy.com/code-of-conduct

OpenStudy (karatechopper):

But..forget it. now i am done. here.